if the perpendicular distance between the man's quadriceps tendon and his knee joint center of rotation remained the same (5cm), how much force would his quadriceps muscle have had to produce at this point?"
1
Expert's answer
2020-04-08T13:32:04-0400

D=5cm

"F=(mg)*4"

If it lifts a mass of 10kgs

F=10*10*4=400N